How many of the numbers from 10 through 83 have the sum of their digits equal to a perfect​ square?
Answer by ikleyn(52835)

.
1.  Of numbers from 10 to 19, the sums of their digits are the integers from 1 to 10.
        Of them, there are 3 numbers with the sums of digits 4 and 9.   (3)


2.  Of numbers from 20 to 29, the sums of their digits are the integers from 2 to 11.
        Of them, there are 2 numbers with the sums of digits 4 and 9.   (2)


3.  Of numbers from 30 to 39, the sums of their digits are the integers from 3 to 12.
        Of them, there are 2 numbers with the sums of digits 4 and 9.   (2)


4.  Of numbers from 40 to 49, the sums of their digits are the integers from 4 to 13.
        Of them, there are 2 numbers with the sums of digits 4 and 9.   (2)


5.  Of numbers from 50 to 59, the sums of their digits are the integers from 5 to 13.
        Of them, there is 1 number with the sum of digits 9.            (1)


6.  Of numbers from 60 to 69, the sums of their digits are the integers from 6 to 14.
        Of them, there is 1 number with the sum of digits 9.            (1)


7.  Of numbers from 70 to 79, the sums of their digits are the integers from 7 to 14.
        Of them, there is 1 number with the sum of digits 9.            (1)


8.  Of numbers from 80 to 83, the sums of their digits are the integers from 8 to 11.
        Of them, there is 1 number with the sum of digits 9.            (1)


9.  Now compute the sum of the numbers in parentheses

    3 + 3*2 + 4*1 = 3 + 6 + 4 = 13


    to get the answer to the problem's question.

Answer.   Among the numbers from  10  through  83,  there are  13  numbers that have the sum of their digits to a perfect square.